OpenStudy (anonymous):

if the area of a base of a rectangular solid is tripled, what is the percent increase in the volume?

OpenStudy (anonymous):

V=area of a base *h => V=3area of base*h

OpenStudy (anonymous):

Can you simplify it a bit more? The answer is supposed to be 200%. But i am getting 300%

OpenStudy (anonymous):

Let x be the area of the base in a rectangular solid. \[V_{100percent}=x\times h\] \[\large V_{?}=3x\times h\] Cross multiply. \[?=\frac{300(x \times h)}{x\times h}\] \[?=300percent\] 300 percent - 100 percent=?

OpenStudy (anonymous):

The question is asking for the percent "increase" of the volume.

OpenStudy (campbell_st):

ok... use an example to solve the problem a rectangular prism, base area = 1 cm^2 ahd height of 1 cm V = 1 x 1 V = 1cm^3 so now triple the area A = 3cm^2 same height V = 3 x 1 V = 3cm^3 so the volume when the base area is tripled is 3 times as large.. the volume is alos 3 times as large so the volume has increase 2cm^3 to the percentage increase is increase/original x 100 or Percentage Increase = 2/1 x 100 = 200%